OpenStudy (anonymous):

I have no idea how to start! 2) The mean of a set of six scores is 20 2a) What is the sum of these six scores? 2b) A seventh score is added and the mean of the seven scores is 22. What is the sum of the seven scores? 2c) What is the seventh score?

jimthompson5910 (jim_thompson5910):

Let S be the sum of the six scores So the mean is equal to the sum (S) divided by the number of scores (6), which is 20, so S/6 = 20 Solve for S

OpenStudy (anonymous):

S = 120?

jimthompson5910 (jim_thompson5910):

good, that's the answer to 2a

jimthompson5910 (jim_thompson5910):

in 2b, it becomes (S+x)/7 = 22 (120+x)/7 = 22 Multiply both sides by 7 to find 120+x, which is the new sum

jimthompson5910 (jim_thompson5910):

Then solve for x to find that 7th score

OpenStudy (anonymous):

Ok. I'll do that now

OpenStudy (anonymous):

34

jimthompson5910 (jim_thompson5910):

you got it, the 7th score is 34

OpenStudy (anonymous):

Thanks so much @jim_thompson5910 !

jimthompson5910 (jim_thompson5910):

yw